OpenStudy (anonymous):

The zeros for (squareroot of x-1) + (squareroot of x + 4) -5 is (are)? Written better below

OpenStudy (anonymous):

The zeros for \[\sqrt{x-1}+\sqrt{x+4} -5\] is (are)?

jhonyy9 (jhonyy9):

your opinion ?

OpenStudy (anonymous):

(a) x = 5 (b) x = 10; 17 (c) x = 1 (d) x = 4

jhonyy9 (jhonyy9):

ok but how you will get these roots of this function so yes this after you make it equale with zero ?

jhonyy9 (jhonyy9):

came on i wann to help you

jhonyy9 (jhonyy9):

how do you think what we need to doing in the first step ?

OpenStudy (anonymous):

im not sure

jhonyy9 (jhonyy9):

hope so much that you have learned in your schoole ,or not ?

OpenStudy (anonymous):

i want to square everything but i feel like thats wrong

OpenStudy (anonymous):

can you please walk me through it

jhonyy9 (jhonyy9):

yes this will be the first step because you know it right that when there is squareroot in afunction first we need to elimine it by squaring it

jhonyy9 (jhonyy9):

so there is sqrt(x-1) +sqrt(x+4) -5 =0

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

sorry my internet conection is wrong

jhonyy9 (jhonyy9):

so in the first step add to both sides 5

jhonyy9 (jhonyy9):

so what will get

jhonyy9 (jhonyy9):

sqrt(x-1) +sqrt(x+4) =5

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

now you need to square both sides ,ok ?

OpenStudy (anonymous):

i got that far and it was x-1+x+4=25

jhonyy9 (jhonyy9):

no you need to square both sides like a=b so but in this your case a =((sqrt(x-1) +sqrt(x+4)) ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than you need to us formula (a+b)^2 = ? hope that you know it

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

\[(\sqrt{x-1}+\sqrt{x+4})^{2}=25\]

OpenStudy (anonymous):

IS that right?

jhonyy9 (jhonyy9):

yes continue it

jhonyy9 (jhonyy9):

do you know formula ?

OpenStudy (anonymous):

no sadly

jhonyy9 (jhonyy9):

so (a+b)^2 =a^2 +2ab +b^2

jhonyy9 (jhonyy9):

a=sqrt(x-1) b=sqrt(x+4)

OpenStudy (anonymous):

ok so its \[x ^{2}+3x-4\]

OpenStudy (anonymous):

=25

jhonyy9 (jhonyy9):

so i dont know it just if you write it here all calcules

jhonyy9 (jhonyy9):

sqrt(x-1) +sqrt(x+4)=5 (x-1) +2sqrt((x-1)(x+4)) +(x+4)=25 2sqrt((x-1)(x+4))=25-(x-1)-(x+4) 2sqrt((x-1)(x+4))=25-x+1-x-4 2sqrt((x-1)(x+4))=22-2x 2sqrt((x-1)(x+4))=2(11-x) sqrt((x-1)(x+4))=11-x (x-1)(x+4)=121-22x +x^2 x^2 +3x -4=x^2 -22x +121 25x =125 x=5 ok ?

OpenStudy (anonymous):

ok thank you

jhonyy9 (jhonyy9):

check it please step by step for you can understanding all right

jhonyy9 (jhonyy9):

was my pleasure good luck bye

OpenStudy (anonymous):

\[\sqrt{x-1}+\sqrt{x+4}=5\]

OpenStudy (anonymous):

\[(x-1)+2(\sqrt{(x-1)(x+4)}+(x+4)=25\]

jhonyy9 (jhonyy9):

yes but next the 2 without paranthese ok ?

OpenStudy (anonymous):

ok